Let m denote the no. of ways in which 4 different books are distributed among 10 persons, each receiving none or one only and let n denote the no. of ways of distribution if the books are all alike. Then:


(A)   m = 4n


(B)   n = 4m


(C)   m = 24n


(D)   none


Grade:Upto college level

3 Answers

Roshan Mohanty
10 years ago

m=10C4  *4!
and n=10C4
so m=n*4!
       =n*24
ans is C 

Vikas TU
10 years ago

different books: m =10P4
identical books: n =10C4

m/n = 4! = 24
-----------------

 

(C)   m = 24n

 

plz approve

jaymin prajapati
10 years ago

different books: m =10P4
identical books: n =10C4

m/n = 4! = 24